# On-device Anomaly Detection in Conveyor Belt Operations

### v2

@@ Abstract
Conveyor belts are crucial in mining operations by enabling the continuous and efficient movement of bulk materials over long distances, which directly impacts productivity. While detecting anomalies in specific conveyor belt components has been widely studied, identifying the root causes of these failures, such as changing production conditions and operator errors, remains critical. Continuous monitoring of mining conveyor belt work cycles is still at an early stage and requires robust solutions. Recently, an anomaly detection method for duty cycle operations of a mining conveyor belt has been proposed. Based on its limited performance and unevaluated long-term proper operation, this study proposes two novel methods for classifying normal and abnormal duty cycles. The proposed approaches are pattern recognition systems that make use of threshold-based duty-cycle detection mechanisms, manually extracted features, pattern-matching, and supervised tiny machine learning models. The explored low-computational models include decision tree, random forest, extra trees, extreme gradient boosting, Gaussian naive Bayes, and multi-layer perceptron. A comprehensive evaluation of the former and proposed approaches is carried out on two datasets. Both proposed methods outperform the former method, with the best-performing approach being dataset-dependent. The heuristic rule-based approach achieves the highest performance in the same dataset used for algorithm training, with 97.3% for normal cycles and 80.2% for abnormal cycles. The ML-based approach performs better on a dataset including the effects of machine aging, scoring 91.3% for normal cycles and 67.9% for abnormal cycles. Implemented on two low-power microcontrollers, the methods demonstrate efficient, real-time operation with energy consumption of 13.3 and 20.6 $\mu$J during inference. These results offer valuable insights for detecting mechanical failure sources, supporting targeted preventive maintenance, and optimizing production cycles.
@@ I Introduction
This work is an extension of the former study by comparing the existing approach with two alternative approaches. The proposed methods are lightweight pattern-recognition systems that leverage manually extracted feature sets and a TinyML classifier [18] to identify the internal operation modes. In the first approach, pattern-matching rules analyze the sequence of internal operation modes to classify duty cycles as normal or abnormal. The second approach replaces these heuristic rules with another TinyML classifier trained to detect anomalous cycles while accounting for potential misclassifications in internal operation modes. Six supervised classifiers are evaluated: decision tree (DT), random forest (RF), extra trees (ET) [19], extreme gradient boosting (XGB) [20], Gaussian naive Bayes (NB), and multi-layer perceptron (MLP) neural network [21]. The best-performing models are deployed on two low-power MCUs for real-time anomaly detection.
@@ II Conveyor Belt Operation
A hydraulic conveyor belt system operates by harnessing fluid power to drive the movement of the belt through hydraulic components. It is a highly efficient continuous transport mechanism, capable of moving heavy loads over long distances. In this system, hydraulic charge pumps pressurize fluid, which is directed through control valves. These valves regulate the flow and direction of oil to drive the hydraulic motors attached to the drive pulleys or rollers of the conveyor belt. The resulting rotational motion propels the belt, facilitating the seamless transportation of materials along the designated path. The control system provides flexibility in adjusting the speed and direction of the belt. It also offers precise control over its operation. Additionally, sensors are connected to the control system. These sensors continuously track and log key hydraulic parameters.
@@
  • Idle: the charge pumps are running, but the belt remains stationary.
  • Operational: the charge pumps and belt are active, but no load is transported.
  • Active: the charge pumps and belt operate, and the belt is transporting material.
@@ III-A Datasets
This study utilized two datasets, each derived from parameters monitored by the control system of a conveyor belt. The conveyor control system maintains a fixed belt speed regardless of material load. Variations in material load alter the required motor torque, causing proportional changes in hydraulic pressure. Monitoring these pressure fluctuations enables real-time assessment of conveyor workload and system stress. Two pressure sensors recorded the high- and low-pressure of the hydraulic motor in bar. Additionally, an incremental encoder with a pulsed output measured the number of motor revolutions per minute (rpm). These sensors were integrated into the control system through a 4-20 mA interface. The collected sensor data underwent preprocessing to achieve a sampling rate of one measurement per minute. These values were retrieved from the control system’s user interface. For each sensor, the one-minute interval values represent the average of all readings collected at the sensor’s nominal sampling frequency.
@@
The first dataset (DS1) comprises data from four months representing all four seasons: June 2021, October 2021, January 2022, and April 2022. This temporal coverage captured a broad temperature range (-26.1 to 26.0°C), significantly affecting oil viscosity and consequently the pressure levels. The second dataset (DS2) was collected from the same conveyor belt during June, August, October, and December 2023, with temperatures ranging from -29.0 to 26.0°C. Figure 1 shows the distribution of the three sensed variables in DS1 and DS2: high-pressure, low-pressure, and speed. Whisker values in the box plots represent the minimum and maximum values. Both datasets show similar spreads for most high-pressure values, though DS2 covers a wider range. DS1’s low-pressure values cover a greater range than DS2’s, with notable quartile differences. This difference could be caused by machine wear and maintenance. Speed values cluster around 0 and 45 rpm in both datasets when the conveyor belt is stopped or moving, respectively. Variations between these values occur during belt acceleration or deceleration.
TABLE I: Operation modes samples in DS1.
@@
TABLE II: Duty cycle samples distribution in DS1 and DS2.
Dataset Samples (percentage) Class
DS1 496 (82.3%) Normal
107 (17.7%) Abnormal
DS2 751 (79.5%) Normal
194 (20.5%) Abnormal
DS1 contains approximately 170,000 operation mode samples in 1-minute intervals, labeled shown in Table I. Table II presents the number of normal and abnormal duty cycles in each dataset. DS1 was used to develop and evaluate the proposed approaches, while DS2 was used to assess the generalization and robustness of the approaches in a different dataset.
Hydraulic system experts analyzed the waveform patterns using the Imagimob Studio software (Imagimob AB, Stockholm, Sweden) to determine the start and end timestamps, and the ground truth labels for the operation modes and duty cycles.
@@ III-B Algorithms
    Approach-1 proposed by Martinez-Rau et al. [17], consists of four consecutive stages (Fig. 2(a)) and serves as a baseline. The algorithm takes as input the three 1-min sensed variables (motor shaft rotational speed, and the high- and low-pressure of the hydraulic motor), which are used to extract a set of 12 features. Four features consist of the three input values plus the differential pressure between the high- and low-pressure. The remaining eight features are derived from the former four features. These are third- and fifth-order moving averages, which effectively capture fast and slow variations in the input variables. The complete feature set feeds an ML model to classify the operation modes labeled every 1 min. Since variations in the operation modes typically occur at lower frequencies, a third-order median filter smooths the predicted labels to reduce label fragmentation. Finally, transitions in the filtered operation mode labels are analyzed through pattern matching to detect and classify normal and abnormal duty cycles. Therefore, the correct detection and classification of duty cycles depends on the correct prediction of the operation modes.
@@
    Approach-2 offers an alternative method for detecting duty cycles (Fig. 2(b)). Approach-1 relies on accurately classifying operation mode labels to detect and classify duty cycles. In contrast, Approach-2 bypasses operation mode classification and focuses directly on input speed values. Based on the speed values distribution (Fig. 1), an empirical threshold of 5 rpm is chosen to determine the boundaries of the duty cycles. Once these boundaries are identified, the remaining steps in Approach-2 follow a similar process to those in Approach-1.
  3. 3.
@@
The operation mode and the duty-cycle classifiers are evaluated with different combination of supervised ML algorithms: DT, RF, ET, XGB, NB, and MLP.
@@ III-C Experimental Setup
64 MHz 256 kB 1 MB 52 $\mu$A/MHz 3.3 V
RP2040 [23] Dual-core ARM Cortex M0+ Raspberry Pi Pico 133 MHz 256 kB 2 MB $\times$ 75 $\mu$A/MHz 3.3 V
@@ IV Results and Discussion
Approach
Operation mode
classifier
DS1 DS2
Approach-1 NB 89.3 61.6
DT 94.7 97.9
XGB 94.4 96.4
ET 95.0 82.4
MLP 92.5 87.8
RF 94.2 98.0
Approach-2 - 97.6 99.2
Approach-3
@@ IV-A Recognition Performance
The detection and classification performance of the approaches using different ML models is presented in terms of F1-scores, averaged over repeated experiments with varying seeds of initialization. Table IV shows the detection rates for operation cycles across approaches. In Approach-1, correct cycle detection relies on accurate classification of operation modes, yielding consistent performance above 89.3% across all ML models in DS1. In DS2, the detection performance varies across ML models. While RF, DT, and XGB models show improvement over DS1, MLP, ET, and NB experience declines. In contrast, Approach-2 and Approach-3 detect cycles independently of the operation mode classifier, using threshold-based detection (see Fig. 2(b) and Fig. 2(c)). This results in 2.5% to 8.2% improvements in DS1 and 1.3% to 37.6% in DS2 over Approach-1.
@@
In DS1, Approach-1 with the ET classifier achieves the highest performance, with an overall F1-score of 92.4%, distinguishing well between normal (97.1%) and abnormal (71.1%) cycles (Fig. 3a). The performance ranking is ET $>$ MLP $>$ RF $>$ XGB $>$ NB $>$ DT, with overall F1-scores decreasing from 92.4% to 83.9%. In DS2, however, RF outperforms other models with an overall F1-score of 83.7%, attaining 64.0% for abnormal and 89.7% for normal cycles, suggesting superior generalization. Other models show a decline in abnormal cycle classification performance by 7.2-25.9% and 8.0-16.4% for normal cycles.
Approach-2 demonstrates substantial improvements over Approach-1 across both datasets (Fig. 3b). In DS1, Approach-2 maintains a comparable F1-score for normal cycles and shows a 7.5-24.6% improvement in abnormal cycle classification. ET again leads with an overall F1-score of 94.0%, achieving 97.3% for normal and 80.2% for abnormal cycles, improving overall performance by 1.7% and abnormal classification by 9.0% over Approach-1. In DS2, normal cycle classification with Approach-2 models is 0.5-4.1% lower than Approach-1. However, abnormal classification improves by 12.4%, 20.2%, and 35.6% when using the MLP, ET, or NB models, respectively. This boosts the overall F1-scores by 3.9%, 8.9%, and 15.0%. Approach-2 with the ET achieves the highest abnormal cycle F1-score, while RF excels in normal cycle and overall F1-scores.
@@
[t] Approach Quantization resolution F1-score [%] Arduino Nano BLE Sense 33 Lite Raspberry Pi Pico Abnormal Normal Overall Energy consumption [$\mu$J/inference] Flash memory [kB] RAM [kB] Energy consumption [uJ/inference] Flash memory [kB] RAM [kB] Approach-21 32-bit float 64.4 87.8 81.4 18.8 265.9 43.1 82.7 209.5 10.4 8-bit integer 64.9 87.9 81.6 13.7 142.4 43.0 13.3 116.4 10.3 Approach-32 32-bit float 65.5 89.2 83.5 26.5 254.9 47.7 40.9 201.3 15.2 8-bit integer 66.4 89.7 84.1 24.4 242.0 47.4 20.6 191.6 15.0
@@
    Using ET as operation mode classifier.
@@
    Using MLP and ET as operation mode and duty cycle-classifiers, respectively.
@@ IV-B MCU Energy Requirements
Energy consumption calculations focused on data representing duty cycles, ensuring that all system stages are active (Fig. 2), capturing the peak energy consumption. Reported energy consumption reflects the end-to-end computation for each 1-min input data, averaged across DS2. After completing the inference, the idle time can be utilized for additional tasks. These include MCU peripheral management, conveyor control, or putting the MCU into sleep mode. In the 32-bit float version, the Arduino Nano BLE Sense 33 Lite consumes less energy than the Raspberry Pi Pico. This is because it features a floating-point computation unit. In contrast, the Raspberry Pi Pico outperforms the Arduino for 8-bit integer representations. Quantizing Approach-2 reduces program memory usage by 46.4% on the Arduino Nano BLE Sense 33 Lite and 44.4% on the Raspberry Pi Pico, while Approach-3 presents reductions of 5.1% and 4.8%, respectively. Across both MCUs, 32-bit float and 8-bit integer versions required similar amounts of RAM, as both store data in 32-bit registers. We hypothesize that the power consumption of quantized approaches could be further reduced by using ultra-low-power 8-bit MCUs, such as the STM8L series MCUs [27].

### v3

@@ Abstract
Conveyor belts are crucial in mining operations by enabling the continuous and efficient movement of bulk materials over long distances, which directly impacts productivity. While detecting anomalies in specific conveyor belt components has been widely studied, identifying the root causes of these failures, such as changing production conditions and operator errors, remains critical. Continuous monitoring of mining conveyor belt work cycles is still at an early stage and requires robust solutions. Recently, an anomaly detection method for duty cycle operations of a mining conveyor belt has been proposed. Based on its limited performance and unevaluated long-term proper operation, this study proposes two novel methods for classifying normal and abnormal duty cycles. The proposed approaches are pattern recognition systems that make use of threshold-based duty-cycle detection mechanisms, manually extracted features, pattern-matching, and supervised tiny machine learning models. The explored low-computational models include decision tree, random forest, extra trees, extreme gradient boosting, Gaussian naive Bayes, and multi-layer perceptron. A comprehensive evaluation of the former and proposed approaches is carried out on two datasets. Both proposed methods outperform the former method in anomaly detection, with the best-performing approach being dataset-dependent. The heuristic rule-based approach achieves the highest F1-score in the same dataset used for algorithm training, with 97.3% for normal cycles and 80.2% for abnormal cycles. The ML-based approach performs better on a dataset including the effects of machine aging, with an F1-score scoring 91.3% for normal cycles and 67.9% for abnormal cycles. Implemented on two low-power microcontrollers, the methods demonstrate efficient, real-time operation with energy consumption of 13.3 and 20.6 µJ during inference. These results offer valuable insights for detecting mechanical failure sources, supporting targeted preventive maintenance, and optimizing production cycles.
@@
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
This work proposes two alternative lightweight pattern-recognition approaches. The methods leverage manually extracted feature sets and a TinyML classifier [18] to identify the internal operation modes. In the first approach, pattern-matching rules analyze the sequence of internal operation modes to classify duty cycles as normal or abnormal. The second approach replaces these heuristic rules with another TinyML classifier trained to detect anomalous cycles while accounting for potential misclassifications in internal operation modes. Six supervised classifiers are evaluated: decision tree (DT), random forest (RF), extra trees (ET) [19], extreme gradient boosting (XGB) [20], Gaussian naive Bayes (NB), and multi-layer perceptron (MLP) neural network [21]. The best-performing models are deployed on two low-power MCUs for real-time anomaly detection.
@@ II Conveyor Belt Operation
A hydraulic conveyor belt system operates by harnessing fluid power to drive the movement of the belt through hydraulic components. It is a highly efficient continuous transport mechanism, capable of moving heavy loads over long distances. In this system, an electric motor, powered by a starter or frequency converter, activates a hydraulic pump. The pump pressurizes fluid, which is first filtered to remove contaminants, ensuring system longevity, and then circulated through a cooler to maintain optimal operating temperatures. The pressurized fluid is directed through control valves, which regulate the flow and direction of oil to drive the hydraulic motors attached to the drive pulleys or rollers of the conveyor belt. These motors convert hydraulic energy into rotational motion, propelling the belt while the hydraulic fluid returns to a storage tank for recirculation.
Operators control and monitor the control system via a control panel unit, connected to the electric motors, pumps, and valves. The unit is equipped with sensors that continuously track and log key parameters. Operators use this information to adjust speed, direction, or load distribution of the conveyor belt, ensuring precise and efficient operation.
@@
  • Idle: the hydraulic pumps are running, but the belt remains stationary.
  • Operational: the hydraulic pumps and belt are active, but no load is transported.
  • Active: the hydraulic pumps and belt operate, and the belt is transporting material.
@@ III-A Datasets
The conveyor belt system used in this study operates under a control scheme that tries to maintain a fixed belt speed irrespective of the material load. As the load on the conveyor varies, the torque demand on the hydraulic motor changes proportionally, resulting in corresponding fluctuations in hydraulic pressure. These pressure variations serve as a proxy for assessing real-time system workload and mechanical stress. The system is instrumented with three sensors: an incremental encoder, and two hydraulic pressure transducers. The encoder, mounted on the hydraulic motor shaft, measures the motor speed in revolutions per minute (rpm) at high resolution (i.e., thousands of pulses per revolution). The high-pressure transducer is positioned at the pump outlet or hydraulic motor inlet and measures the pressure of the fluid that drives the motor in bar. This pressure directly reflects the load conditions, as higher loads require greater pressure to maintain constant belt speed. The low-pressure transducer, placed at the pump inlet or on the motor return line to the tank, monitors return-line pressure in bar to ensure unrestricted fluid flow back to the reservoir. Elevated low-pressure values may indicate issues such as filter clogging or cooler obstruction. Both pressure sensors communicate with the control system via 4–20 mA analog signals, while the encoder is connected directly. Sensor data are sampled at 10 Hz by the control system for setpoint updates and local logging. For long-term monitoring and storage efficiency, the system averages each sensor’s readings over one-minute intervals and transmits these values to a secure central server. This study utilized two datasets retrieved through a web-based interface, comprising these one-minute interval records and enabling analysis of the conveyor belt’s operational parameters over time.
@@
The first dataset (DS1) was collected over four months (June 2021, October 2021, January 2022, and April 2022), ensuring coverage of all four seasons and a broad temperature range (-26.1 to 26.0 °C). The second dataset (DS2) was acquired from the same conveyor belt during June, August, October, and December 2023, with temperatures ranging from -29.0 to 26.0 °C. These temperature variations significantly influenced oil viscosity, thereby affecting system pressure levels. Each dataset comprises approximately 170,000 samples, where each sample represents one-minute averaged measurements of speed, high-pressure, and low-pressure values. Figure 1 shows the distribution of the three sensed variables in DS1 and DS2: high-pressure, low-pressure, and speed. Whisker values in the box plots represent the minimum and maximum values. Both datasets exhibit similar spreads for most high-pressure values, although DS2 spans a wider range. DS1’s low-pressure values span a greater range than those of DS2, with notable quartile differences. This difference could be caused by machine wear and maintenance. Speed values cluster around 0 and 45 rpm in both datasets when the conveyor belt is stopped or moving, respectively. Variations between these values occur during belt acceleration or deceleration.
Hydraulic system experts analyzed the waveform patterns of the data using the DEEPCRAFT™ Studio software (Infinion Technologies AG, Neubiberg, Germany) to determine the start and end timestamps, and the ground truth labels for the operation modes and duty cycles, described in Section II. Table I presents the number of normal and abnormal duty cycles in each dataset. Due to the extensive time consumption of the labeling task, operation mode labels are available exclusively on DS1. These labels were 1-minute segmented to match the three sensed variables, and their distribution is presented in Table II. DS1 was used to develop and evaluate the proposed approaches, while DS2 was used to assess the generalization and robustness of the approaches in a different dataset.
TABLE I: Duty cycle samples distribution in DS1 and DS2.
Dataset Samples (percentage) Class
DS1 496 (82.3%) Normal
107 (17.7%) Abnormal
DS2 751 (79.5%) Normal
194 (20.5%) Abnormal
TABLE II: Operation modes samples in DS1.
@@ III-B Algorithms
    Approach-1 proposed by Martinez-Rau et al. [17], consists of four consecutive stages (Fig. 2(a)) and serves as a baseline. The algorithm takes as input the three 1-min sensed variables (motor shaft rotational speed, and the high- and low-pressure of the hydraulic motor), which are used to extract a set of 12 features. Four features consist of the three input values plus the differential pressure between the high- and low-pressure. Differential pressure determines the motor torque output. The remaining eight features are derived from the former four features. These are third- and fifth-order moving averages, which effectively capture fast and slow variations in the input variables. The complete feature set feeds an ML model to classify the operation modes labeled every 1 min. Since variations in the operation modes typically occur at lower frequencies, a third-order median filter smooths the predicted labels to reduce label fragmentation. Finally, transitions in the filtered operation mode labels are analyzed through pattern matching to detect and classify normal and abnormal duty cycles. The main drawback of this method is that duty-cycle detection depends on correct operation mode classification and good temporal alignment of the pattern matching analyzer. Approach-2 offers an alternative solution by decoupling the detection and classification tasks. Instead of relying solely on operation mode classification, this method first analyzes input speed values to identify the start and end of a duty cycle. Based on the speed values distribution (Fig. 1), an empirical threshold of 5 rpm is chosen to determine the boundaries of the duty cycles. Once these boundaries are identified, the remaining steps in Approach-2 follow a similar process to those in Approach-1.
@@
The operation mode and the duty-cycle classifiers are evaluated with different combinations of supervised ML algorithms: DT, RF, ET, XGB, NB, and MLP.
@@ III-C Experimental Setup
64 MHz 256 kB 1 MB 52 µA/MHz 3.3 V
RP2040 [23] Dual-core ARM Cortex M0+ Raspberry Pi Pico 133 MHz 256 kB 2 MB $\times$ 75 µA/MHz 3.3 V
@@ IV Results and Discussion
[t] Approach Detection method DS1 DS2 Approach-1* NB 89.3 61.6 DT 94.7 97.9 XGB 94.4 96.4 ET 95.0 82.4 MLP 92.5 87.8 RF 94.2 98.0 Approach-2 TBD** 97.6 99.2 Approach-3
    Duty-cycle detection depends on the operation mode classifier.
    Threshold-based detection (TBD) independent of the classifier.
@@ IV-A Recognition Performance
The detection and classification performance of the approaches using different ML models is presented in terms of F1-scores, averaged over repeated experiments with varying seeds of initialization. Table IV shows the detection rates for operation cycles across approaches. In Approach-1, correct cycle detection relies on accurate classification of operation modes, yielding consistent performance above 89.3% across all ML models in DS1. In DS2, the detection performance varies across ML models. While RF, DT, and XGB models show improvement over DS1, MLP, ET, and NB experience declines. In contrast, Approach-2 and Approach-3 detect cycles independently of the operation mode classifier, using threshold-based detection (TBD) (see Fig. 2(b) and Fig. 2(c)). This results in 2.5% to 8.2% improvements in DS1 and 1.3% to 37.6% in DS2 over Approach-1.
@@
In DS1, Approach-1 with the ET classifier achieves the highest performance, with an overall F1-score of 92.4%, distinguishing well between normal (97.1%) and abnormal (71.1%) cycles (Fig. 3a). The performance ranking is ET >> MLP >> RF >> XGB >> NB >> DT, with overall F1-scores decreasing from 92.4% to 83.9%. In DS2, however, RF outperforms other models with an overall F1-score of 83.7%, attaining 64.0% for abnormal and 89.7% for normal cycles, suggesting superior generalization. Other models show a decline in abnormal cycle classification performance by 7.2-25.9% and 8.0-16.4% for normal cycles.
Approach-2 (Fig. 3b) and Approach-3 (Fig. 4 and Fig. 5) demonstrate substantial improvements over Approach-1 across both datasets. In DS1, Approach-2 maintains a comparable F1-score for normal cycles and shows a 7.5-24.6% improvement in abnormal cycle classification. ET again leads with an overall F1-score of 94.0%, achieving 97.3% for normal and 80.2% for abnormal cycles, improving overall performance by 1.7% and abnormal classification by 9.0% over Approach-1. In DS2, normal cycle classification with Approach-2 models is 0.5-4.1% lower than Approach-1. However, abnormal classification improves by 12.4%, 20.2%, and 35.6% when using the MLP, ET, or NB models, respectively. This boosts the overall F1-scores by 3.9%, 8.9%, and 15.0%. Approach-2 with the ET achieves the highest abnormal cycle F1-score, while RF excels in normal cycle and overall F1-scores. These results confirm that the pattern matching technique for classifying duty-cycles is effective, but is not suitable for their detection.
@@
[t] Approach Quantization resolution F1-score [%] Arduino Nano BLE Sense 33 Lite Raspberry Pi Pico Abnormal Normal Overall Energy consumption [µJ/inference] Flash memory [kB] RAM [kB] Energy consumption [uJ/inference] Flash memory [kB] RAM [kB] Approach-21 32-bit float 64.4 87.8 81.4 18.8 265.9 43.1 82.7 209.5 10.4 8-bit integer 64.9 87.9 81.6 13.7 142.4 43.0 13.3 116.4 10.3 Approach-32 32-bit float 65.5 89.2 83.5 26.5 254.9 47.7 40.9 201.3 15.2 8-bit integer 66.4 89.7 84.1 24.4 242.0 47.4 20.6 191.6 15.0
@@
    Using ET with 50 trees as operation mode classifier.
@@
    Using MLP with 12 neurons in the hidden layer as operation mode classifier and ET with 50 trees as duty cycle classifier.
@@ IV-B MCU Energy Requirements
Energy consumption calculations focused on data representing duty cycles, ensuring that all system stages are active (Fig. 2), capturing the peak energy consumption. Reported energy consumption reflects the end-to-end computation for each 1-min input data, averaged across DS2. After completing the inference, the idle time can be utilized for additional tasks. These include MCU peripheral management, conveyor control, or putting the MCU into sleep mode.
The Arduino Nano BLE Sense 33 Lite, which features a floating-point computation unit (FPU), consumes less energy than the Raspberry Pi Pico when running the 32-bit float versions of the approaches. The effect of quantization to 8-bit integer on the Arduino Nano BLE Sense 33 Lite reduces energy consumption by 27.1% for Approach-2 and 7.9% for Approach-3. This reduction occurs because operations performed by the FPU in the non-quantized versions are executed by the CPU in the quantized versions. The impact of quantization is more pronounced on the Raspberry Pi Pico, decreasing energy usage by 83.9% for Approach-2 and 49.6% for Approach-3.
The cycle classifier contributes minimally to overall energy consumption since it performs only a single inference at the end of each cycle. Thus, the key difference in energy usage between Approach-2 and Approach-3 arises from their distinct operational mode classifiers. Notably, Approach-3 consumes more energy than Approach-2 on the Arduino Nano BLE Sense 33 Lite. This is because the ET classifier (Approach-2) primarily relies on logical comparisons, which can be executed in a single clock cycle on the FPU, whereas the MLP classifier (Approach-3) requires more complex operations related to activation functions. Conversely, the opposite behavior is observed for the floating-point versions of the approaches on the Raspberry Pi Pico. For this MCU, the C code compiler inserts additional code to emulate floating-point operations using integer arithmetic, resulting in more overhead for ET than for the MLP.
Quantizing Approach-2 reduces program memory usage by 46.4% on the Arduino Nano BLE Sense 33 Lite and 44.4% on the Raspberry Pi Pico, while Approach-3 presents reductions of 5.1% and 4.8%, respectively. Across both MCUs, 32-bit float and 8-bit integer versions required similar amounts of RAM, as both store data in 32-bit registers. The low resource usage of the deployed methods in the MCUs suggests their compatibility with other general-purpose MCUs. Building on this efficiency, we hypothesize that the power consumption of quantized approaches could be further reduced by using ultra-low-power 8-bit MCUs, such as the STM8L series MCUs [27].